what is the focal diameter of: -48(y-(-9))=(x-(-8))^2
Answer by lwsshak3(11628)

what is the focal diameter of: -48(y-(-9))=(x-(-8))^2
**
'
By focal diameter I assume it is another name for focal width=4p.
..
-48(y-(-9))=(x-(-8))^2
-48(y+9)=(x+8)^2
(x+8)^2=-48(y+9)
This is an equation of a parabola of the standard form: (x-h)^2=4p(y-k), with (h,k) being the (x,y) coordinates of the vertex.
For given equation:
Vertex: (-8,-9)
Parabola opens downwards so parabola has a maximum
Focal width=4p=48